OpenStudy (anonymous):

4(x-b)+x PLEASE HELP!

OpenStudy (bibby):

Use the distributive property (a(b-c) = a*b - a*c) 4x-4b+x and combine like terms. What do you get next?

OpenStudy (anonymous):

No idea.

OpenStudy (bibby):

it'll help if you rearrange the terms. 4x-4b+x = 4x+x-4b = (4x+x)-4b do you know what 4x + x is?

OpenStudy (anonymous):

Not a clue, math isn't't my strong suit.

OpenStudy (anonymous):

isn't *

OpenStudy (bibby):

Well, here's some basic rules for adding variables. x+x = 2x 3x+2x = 5x 2x-x = x x*x = x^2 (x squared) you can imagine it as 2 blobs of WHATEVER. And you're adding and subtracting those blobs.